Version 1 Algebra 2 paces provided. (2 points) A standard deck of playing cards is comprised of 52 cards, Ace through King for each of the four suit pattens. Kings, Queens, and Jacks are considered face-cards. Using complements, find the probability of not choosing a face-card from a standard deck of cards. 8
Answer #1

Out of 52 cards, 12 are face cards. So the probability of choosing a face card is

P(face card) = 12 /52

By the complement rule, the probability of not choosing a face card is

P(not face card) = 1 - P(face card) = 1 - 12/52 = 40/52 = 0.7692

Answer: 10/13 or 0.7692